# Emotion-LLaMA: Multimodal Emotion Recognition and Reasoning with Instruction Tuning

### v1

@@ 4.6 Ablation Evaluation
We conducted a series of ablation experiments on the MER2023-SEMI  [43] dataset to explore the effectiveness of each component of the proposed Emotion-LLaMA.
Effect of Sample Selection Strategy. To better understand the effect of the sample selection strategy, we compared our strategy against traditional semi-supervised approaches, as shown in Table 4.6. Due to the limited size of the MER2023 training dataset, which contains only 3,373 samples, pre-training on this dataset can lead to difficulties in fitting models with transformer structures, resulting in a low F1 score of 0.7977. We then compared traditional semi-supervised approaches, which involve assigning pseudo-labels to all unlabeled samples or selectively giving pseudo-labels to those samples that exhibit high softmax scores during inference. There are 73,148 and 36,490 samples selected by these two strategies, respectively. During the pre-training phase, a substantial increase in data volume can significantly enhance performance, even introducing considerable noise. Ultimately, the model tuning on our automatically annotated MERR dataset achieves the best model performance. This demonstrates the effectiveness and robustness of the proposed Emotion-LLaMA in leveraging large-scale, diverse datasets for improved emotion recognition and reasoning.
Table 5: Ablation Study Results for Different Encoders
@@
Table 6: Ablation Study Results for Different Training Strategies. “Standard Training” involves training solely with the dataset. Both “High-Confidence Training” and “Pseudo-Label Training” are semi-supervised learning strategies; the former uses only high-confidence pseudo-labeled data, while the latter utilizes all available data. “Instruction Tuning” employs annotated data.
@@
5 Conclusion
In this paper, we have introduced a multimodal large language model, Emotion-LLaMA, designed to recognize and understand human emotional states in the real world. Initially, we employed open-source tools for automated selection and annotation of the Multimodal Emotion Recognition and Reasoning (MERR) dataset, on which we pre-trained our model. Subsequently, we performed instruction-tuning on datasets such as MER2023 and DFEW, achieving state-of-the-art (SOTA) scores. We then compared the predictive outcomes of Emotion-LLaMA with other MLLMs, demonstrating our model’s robust generalization capabilities in emotion recognition and reasoning. The experimental results and qualitative analysis presented in this paper highlight the effectiveness of our approach in capturing and interpreting emotional cues from multiple modalities, leading to more accurate and contextually relevant emotion recognition. We believe that Emotion-LLaMA has the potential to advance the field of multimodal emotion understanding and pave the way for more empathetic and emotionally intelligent AI systems.
@@ A.1 Emotion Categories and Annotations
To accurately categorize these challenging emotions, the MERR dataset relies on rich multimodal descriptions that provide a comprehensive understanding of the emotional state and its context. These descriptions go beyond simple categorical labels and offer detailed insights into the facial expressions, body language, vocal cues, and situational factors that contribute to the emotional interpretation. Table 9 presents a template for the multimodal descriptions used in MERR, showcasing the different components that are captured for each sample. The descriptions include a visual expression component that focuses on the specific facial movements and action units associated with the emotion, a visual objective component that describes the overall scene and context, an audio tone component that captures the vocal cues and intonation, and a textual component that provides the transcribed speech or dialogue. To further illustrate the value of these detailed annotations, Tables 10 and  11 provide specific examples of annotated samples for doubt and contempt, respectively.
@@ A.3 Instruction Collection and Multimodal Annotation
Visual Expression Description. In videos, natural actions such as blinking and speaking can lead to different combinations of Action Units (AUs) being extracted from various frames. To accurately represent the current emotion, it is crucial to determine the most relevant AUs. As illustrated in Figure 6, our approach involves analyzing the amplitude values of the AUs to identify the "emotional peak frame", which captures the most intense emotional expression.
@@
    Identify the most frequently occurring AUs across all frames in the video segment.
@@
Visual Objective Description. To provide a comprehensive understanding of the emotional context in each video, we utilize the MiniGPT-v2 [8] model to generate descriptions of the visual scene. By inputting the complete emotional peak frame, which captures the most intense emotional expression, MiniGPT-v2 can analyze and describe various aspects of the video, such as the environment, character actions, and object interactions. These visual objective descriptions offer valuable insights into the situational context surrounding the emotional expression. For example, if a character is shown in a dimly lit room with a concerned facial expression, the model might generate a description like "The scene takes place in a dark, shadowy room. The character appears to be sitting alone, with a worried look on their face, fidgeting with their hands." This description provides additional information about the setting and the character’s body language, which can help in interpreting their emotional state. Moreover, the model can also identify and describe relevant objects or elements in the scene that may contribute to the emotional context. For instance, if a character is holding a letter and appears upset, the model might mention the presence of the letter in its description, suggesting that the content of the letter could be related to the character’s emotional response.
@@
Audio Tone Description. In addition to visual cues, audio plays a crucial role in conveying emotional information. The tone, intonation, and prosodic features of a speaker’s voice can provide valuable insights into their emotional state, often revealing subtle nuances that may not be apparent from visual cues alone. To capture these audio cues, we employ the Qwen-Audio [18] model, which is specifically designed to analyze and describe the emotional content of speech. By processing the audio track of each video segment, it can generate detailed descriptions of the speaker’s tone and intonation. For example, if a character is speaking with a trembling voice and a high pitch, the model might generate a description like "The speaker’s voice is shaky and high-pitched, indicating a sense of fear or anxiety. There are noticeable pauses and hesitations in their speech, suggesting uncertainty or distress." This description captures the emotional nuances conveyed through the speaker’s vocal delivery, providing additional context for understanding their emotional state. Moreover, Qwen-Audio can also identify and describe other relevant audio cues, such as sighs, laughter, or changes in speaking rate, which can further contribute to the emotional interpretation. For instance, if a character is speaking rapidly and with a sarcastic tone, the model might generate a description like "The speaker’s voice is fast-paced and has a sharp, biting tone, conveying a sense of anger or frustration. There are occasional scoffs and short, humorless laughs, indicating a mocking or dismissive attitude."
@@
Multimodal Description. To generate initial coarse-grained emotional descriptions, we concatenate the information obtained from all modalities (visual expression, visual context, audio tone, and textual content). These surface-level descriptions for the 28,618 pseudo-labeled samples are used in the first stage of pre-training to help the model align emotional features with the textual semantic space. We input all the collected emotional clues into the LLaMA-3 model for further refinement. LLaMA-3 sifts through the clues, identifies the most relevant ones, and combines them to generate a comprehensive emotional description. This process helps to filter out any erroneous or contradictory descriptions that may have been present in the initial set of emotional clues. Additionally, we remove duplicate or overabundant samples to ensure a balanced and diverse dataset. Through these refinement processes, the final MERR dataset contains 4,487 samples, each accompanied by a detailed multimodal description. Table 9 presents an example of the annotation format used for each sample in the dataset. By collecting and integrating instructions and descriptions from multiple modalities, we have created a rich and informative dataset that captures the complexities of emotional expressions in real-world scenarios. The multimodal annotations in MERR enable models to learn more comprehensive and nuanced representations of emotions, leading to improved performance on emotion recognition and reasoning tasks.
@@ A.4 Data Statistics and Comparisons
Comparison with Related Datasets. To highlight the unique features and contributions of the MERR dataset, we compare it with several related datasets in Table 12. The MER2023 [43] and DFEW [33] datasets primarily focus on discrete emotion category labels, providing a classification-oriented perspective on emotion recognition. While these datasets are valuable for developing models that can predict specific emotion categories, they lack the detailed descriptions and contextual information necessary for deeper emotion understanding. On the other hand, datasets like EmoSet [70] and EmoVIT [69] offer visual descriptions of emotional expressions, capturing the facial cues and body language associated with different emotions. However, these datasets do not include information from other modalities, such as audio or text, which can provide crucial insights into the emotional context and help disambiguate complex or subtle expressions. The EMER [45] dataset stands out as one of the few existing datasets that contain multimodal emotion descriptions, incorporating information from visual, audio, and textual modalities. However, due to the high cost and effort involved in manual annotation, EMER is limited to only 100 samples, which may not be sufficient for training robust and generalizable emotion recognition models.
@@
Table 10: Example of the emotion category doubt”. The table presents a multimodal data sample associated with the doubt” category, including the visual expression description, audio tone description, and textual content. This example highlights the subtle and complex nature of the “doubt” emotion and the importance of considering multiple modalities for accurate recognition and reasoning.
@@
Table 11: Example of the emotion category contempt”. The table showcases a multimodal data sample corresponding to the contempt” emotion, encompassing the visual expression description, audio tone description, and textual content. This example illustrates the nuanced and challenging characteristics of the “contempt” category, emphasizing the need for a comprehensive multimodal approach to capture its subtleties.
@@ B.3.1 Multi-Task Learning Approach
One of the core aspects of the Emotion-LLaMA model is its ability to simultaneously learn and perform multiple tasks related to emotion understanding. Specifically, the model is trained using a multi-task learning approach, where the emotion recognition and emotion reasoning tasks are learned in parallel. In the emotion recognition task, the model learns to predict the appropriate emotion label for a given multimodal input, such as a video clip accompanied by audio and text. The model is trained to map the input features to one of the predefined emotion categories, such as happy, sad, angry, or neutral. On the other hand, the emotion reasoning task focuses on generating human-interpretable explanations for the predicted emotion labels. Given a multimodal input, the model learns to analyze the various cues and generate a natural language explanation that justifies the predicted emotion based on the available evidence. By training the model to perform both tasks simultaneously, we allow it to share representations and learn complementary skills across tasks. This multi-task learning approach has several benefits. First, it enables the model to develop a more comprehensive understanding of emotions by learning to recognize the emotional state and explain the reasoning behind it. Second, it encourages the model to learn more robust and generalizable representations by leveraging the commonalities and differences between the two tasks. During training, we alternate between the emotion recognition and emotion reasoning tasks, presenting the model with batches of data from each task. The model is optimized using a weighted combination of the task-specific losses, which allows it to balance the learning of both tasks effectively.
@@ B.3.2 Coarse-to-Fine Training Strategy
To facilitate the learning of nuanced and detailed emotional representations, we employ a coarse-to-fine training strategy. This strategy involves two main stages: pre-training on coarse-grained annotations and fine-tuning on fine-grained annotations. In the pre-training stage, the model is trained on the coarse-grained annotations from the MERR dataset. These annotations provide a high-level categorization of emotions, such as positive, negative, or neutral. By training on these coarse-grained labels, the model learns to capture the general emotional tone and develop an initial understanding of the emotional content in the multimodal data. After the pre-training stage, we proceed to the fine-tuning stage, where the model is exposed to the fine-grained annotations from the MERR dataset. These annotations provide more detailed and specific emotion labels, such as happy, sad, angry, worried, or surprised. By fine-tuning the model on these fine-grained labels, we enable it to refine its emotional understanding and capture more nuanced expressions and variations within each emotional category. The coarse-to-fine training strategy offers several advantages. First, it allows the model to gradually learn more complex and detailed emotional representations, starting from a solid foundation of general emotional understanding. Second, it helps to prevent overfitting and improves the model’s generalization ability by providing a hierarchy of emotional labels to learn from. During the fine-tuning stage, we typically use a smaller learning rate and a shorter training duration compared to the pre-training stage. This allows the model to adapt its representations to the fine-grained labels without drastically changing the learned features from the pre-training stage.

### v2

@@ 4.6 Ablation Evaluation
We conducted a series of ablation experiments to explore the effectiveness of each component of the proposed Emotion-LLaMA. More ablation experiments, which examine factors affecting instruction-tuning performance, including data quantity and quality as well as hyperparameters, are presented in Appendix  C.
Table 6: Ablation Study Results for Different Encoders.
@@
In Table 4.6, we present the impact of different instruction data on the instruction-tuning of Emotion-LLaMA. ’Raw’ refers to the direct concatenation of visual and audio descriptions as instructions for training Emotion-LLaMA, which yielded the poorest performance. When trained using the coarse-grained set from the MERR dataset, Emotion-LLaMA achieved scores of 7.41 and 5.56 for clue and label overlap, respectively. This marks an improvement of 1.87 and 1.25 over the ’Raw’ approach, demonstrating that coarse-grained annotations generated by the LLaMA-3 model effectively integrate emotional cues to capture genuine emotional expressions. Notably, further instruction-tuning using the fine-grained set from the MERR dataset resulted in additional gains of 0.42 and 0.69 in clue and label overlap, respectively, indicating that fine-grained annotations offer higher quality data and further enhance the performance of instruction-tuning.
To better understand the effect of the sample selection strategy, we compared our strategy against traditional semi-supervised approaches, as shown in Table 4.6. Due to the limited size of the MER2023 training dataset, which contains only 3,373 samples, pre-training on this dataset can lead to difficulties in fitting models with transformer structures, resulting in a low F1 score of 0.7977. We then compared traditional semi-supervised approaches, which involve assigning pseudo-labels to all unlabeled samples or selectively giving pseudo-labels to those samples that exhibit high softmax scores during inference. There are 73,148 and 36,490 samples selected by these two strategies, respectively. During the pre-training phase, a substantial increase in data volume can significantly enhance performance, even introducing considerable noise. Ultimately, the model tuning on our automatically annotated MERR dataset achieves the best model performance. This demonstrates the effectiveness and robustness of the proposed Emotion-LLaMA in leveraging large-scale, diverse datasets for improved emotion recognition and reasoning.
Table 7: Ablation Study Results for Different Stages of the MERR Dataset Instructions.
Stage Clue Overlap Label Overlap
Raw 5.54 4.31
Coarse-grained 7.41 5.56
Fine-grained 7.83 6.25
Table 8: Ablation Study Results for Different Training Strategies.
@@
5 Ethics and Conclusion
Ethics. All datasets used in this study are governed by signed usage agreements, strictly restricting their use to academic research. The MERR dataset is derived from MER2023 and includes over 70,000 unannotated samples from diverse movies and TV series. We have obtained the necessary End User License Agreements (EULA) and explicit permissions from the original data providers. The open-source MERR dataset contains only emotion description JSON files, intentionally excluding source videos. Researchers must apply directly to the original providers and fully comply with the EULA to access the dataset. We ensure that the MERR dataset exclusively comprises multimodal emotion descriptions without any discriminatory or biased content.
Conclusion. We introduced Emotion-LLaMA, a novel multimodal large language model designed to accurately recognize and interpret human emotions in real-world scenarios. Utilizing robust open-source tools, we automated the selection and annotation of the Multimodal Emotion Recognition and Reasoning (MERR) dataset for pre-training purposes. Instruction-tuning on comprehensive datasets such as MER2023 and DFEW enabled us to achieve state-of-the-art performance metrics. Comparative analyses with other advanced multimodal large language models (MLLMs) demonstrated Emotion-LLaMA’s superior generalization capabilities in emotion recognition and reasoning tasks.
Acknowledgments. This work was partially supported by the National Natural Science Foundation of China (Grant No. 62176165), Stable Support Projects for Shenzhen Higher Education Institutions (Grant No. 20220718110918001), and the Natural Science Foundation of Top Talent at SZTU (Grant No. GDRC202131). Additional support was provided by NSF CISE (Grant No. 1937998), the Air Force Research Laboratory (Grant No. FA8750-19-2-0200), U.S. Department of Commerce (Grant No. 60NANB17D156), IARPA (Grant No. D17PC00340), and DARPA (Grant No. HR00111990063). The research findings and conclusions presented in this paper are solely those of the authors and do not necessarily represent the views of the funding agencies.
@@ A.1 Emotion Categories and Annotations
To accurately categorize these challenging emotions, the MERR dataset relies on rich multimodal descriptions that provide a comprehensive understanding of the emotional state and its context. These descriptions go beyond simple categorical labels and offer detailed insights into the facial expressions, body language, vocal cues, and situational factors that contribute to the emotional interpretation. Table 12 presents a template for the multimodal descriptions used in MERR, showcasing the different components that are captured for each sample. The descriptions include a visual expression component that focuses on the specific facial movements and action units associated with the emotion, a visual objective component that describes the overall scene and context, an audio tone component that captures the vocal cues and intonation, and a textual component that provides the transcribed speech or dialogue. To further illustrate the value of these detailed annotations, Tables 13 and  14 provide specific examples of annotated samples for doubt and contempt, respectively.
@@ A.3 Instruction Collection and Multimodal Annotation
Visual Expression Description. In videos, natural actions such as blinking and speaking can lead to different combinations of Action Units (AUs) being extracted from various frames. To accurately represent the current emotion, it is crucial to determine the most relevant AUs. As illustrated in Figure 6, our approach involves analyzing the amplitude values of the AUs to identify the “emotional peak frame”, which captures the most intense emotional expression.
@@
    Identify the AUs present in all frames of the video segment.
@@
    Determine whether the combinations of these AUs match the pseudo-label.
  4. 4.
@@
Visual Objective Description. To provide a comprehensive understanding of the emotional context in each video, we utilize the MiniGPT-v2 [10] model to generate descriptions of the visual scene. By inputting the complete emotional peak frame, which captures the most intense emotional expression, MiniGPT-v2 can analyze and describe various aspects of the video, such as the environment, character actions, and object interactions. These visual objective descriptions offer valuable insights into the situational context surrounding the emotional expression. For example, if a character is shown in a dimly lit room with a concerned facial expression, the model might generate a description like “The scene takes place in a dark, shadowy room. The character appears to be sitting alone, with a worried look on their face, fidgeting with their hands.” This description provides additional information about the setting and the character’s body language, which can help in interpreting their emotional state. Moreover, the model can also identify and describe relevant objects or elements in the scene that may contribute to the emotional context. For instance, if a character is holding a letter and appears upset, the model might mention the presence of the letter in its description, suggesting that the content of the letter could be related to the character’s emotional response.
@@
Audio Tone Description. In addition to visual cues, audio plays a crucial role in conveying emotional information. The tone, intonation, and prosodic features of a speaker’s voice can provide valuable insights into their emotional state, often revealing subtle nuances that may not be apparent from visual cues alone. To capture these audio cues, we employ the Qwen-Audio [22] model, which is specifically designed to analyze and describe the emotional content of speech. By processing the audio track of each video segment, it can generate detailed descriptions of the speaker’s tone and intonation. For example, if a character is speaking with a trembling voice and a high pitch, the model might generate a description like “The speaker’s voice is shaky and high-pitched, indicating a sense of fear or anxiety. There are noticeable pauses and hesitations in their speech, suggesting uncertainty or distress.” This description captures the emotional nuances conveyed through the speaker’s vocal delivery, providing additional context for understanding their emotional state. Moreover, Qwen-Audio can also identify and describe other relevant audio cues, such as sighs, laughter, or changes in speaking rate, which can further contribute to the emotional interpretation. For instance, if a character is speaking rapidly and laughing, the model might generate a description like “The person in the video speaks with a cheerful tone.”
@@
Multimodal Description. To generate initial coarse-grained emotional descriptions, we concatenate the information obtained from all modalities (visual expression, visual context, audio tone, and textual content). These surface-level descriptions for the 28,618 pseudo-labeled samples are used in the first stage of pre-training to help the model align emotional features with the textual semantic space. We input all the collected emotional clues into the LLaMA-3 model for further refinement. LLaMA-3 sifts through the clues, identifies the most relevant ones, and combines them to generate a comprehensive emotional description. This process helps to filter out any erroneous or contradictory descriptions that may have been present in the initial set of emotional clues. Additionally, we remove duplicate or overabundant samples to ensure a balanced and diverse dataset. Through these refinement processes, the final MERR dataset contains 4,487 samples, each accompanied by a detailed multimodal description. Table 12 presents an example of the annotation format used for each sample in the dataset. Finally, four domain experts manually review the refined descriptions and use a voting process to select fine-grained samples, assessing whether the video descriptions are reasonable and if the emotional reasoning is accurate. By collecting and integrating instructions and descriptions from multiple modalities, we have created a rich and informative dataset that captures the complexities of emotional expressions in real-world scenarios. The multimodal annotations in MERR enable models to learn more comprehensive and nuanced representations of emotions, leading to improved performance on emotion recognition and reasoning tasks.
@@ A.4 Data Statistics and Comparisons
Comparison with Previous Datasets. To highlight the unique features and contributions of the MERR dataset, we compare it with several related datasets in Table 15. The MER2023 [59] and DFEW [45] datasets primarily focus on discrete emotion category labels, providing a classification-oriented perspective on emotion recognition. While these datasets are valuable for developing models that can predict specific emotion categories, they lack the detailed descriptions and contextual information necessary for deeper emotion understanding. On the other hand, datasets like EmoSet [93] and EmoVIT [92] offer visual descriptions of emotional expressions, capturing the facial cues and body language associated with different emotions. However, these datasets do not include information from other modalities, such as audio or text, which can provide crucial insights into the emotional context and help disambiguate complex or subtle expressions. The EMER [62] dataset stands out as one of the few existing datasets that contain multimodal emotion descriptions, incorporating information from visual, audio, and textual modalities. However, due to the high cost and effort involved in manual annotation, EMER is limited to only 100 samples, which may not be sufficient for training robust and generalizable emotion recognition models.
@@
Assessing the Quality of Dataset Instructions. We conducted a human evaluation of the annotation process by randomly selecting 20 video samples from each of the nine emotion categories. We then randomly shuffled the fine-grained annotations. Five volunteers evaluated the consistency and relevance of the video descriptions by scoring them. Each volunteer rated a total of 180 descriptions on a scale from 0 to 5. The evaluation criteria included:
  1. 1.
    Accuracy of the visual modality description.
  2. 2.
    Accuracy of the audio modality description.
  3. 3.
    Accuracy of the textual modality description.
  4. 4.
    Correctness of the reasoning process.
  5. 5.
    Correctness of the reasoning result.
As shown in the Table 9, the average score for the human evaluation of the MERR dataset is 4.258, indicating that the annotations are of high quality and align with real-world logic. Secondly, the "Neutral" category received the lowest score among all categories, suggesting that when a character’s emotion is neutral, the facial and audio cues are weaker, making automatic annotation more challenging. We will address these findings and discuss related limitations in future work. Details of the human evaluation, including the code and assessment results, can be accessed through the code repository.
Table 9: The scores from manual evaluation of the MERR datasets for fine-grained annotation. The scores range from 0 to 5, with the Mean representing the average score across all categories.
Volunteer Angry Happy Surprise Fear Sad Worry Neutral Doubt Contempt Mean
Human-1 4.23 4.33 4.31 4.19 4.27 4.33 4.04 4.55 4.79 4.34
Human-2 4.54 4.78 4.88 4.81 4.60 4.62 4.65 4.55 4.68 4.67
Human-3 3.92 4.11 4.25 4.50 4.00 4.24 3.78 4.27 4.05 4.12
Human-4 3.69 3.94 3.62 3.69 3.93 4.19 3.39 4.00 4.53 3.90
Human-5 3.92 4.72 4.25 4.06 4.00 4.19 4.17 4.32 4.58 4.26
@@
Table 13: Example of the emotion category “doubt”. The table presents a multimodal data sample associated with the “doubt” category, including the visual expression description, audio tone description, and textual content. This example highlights the subtle and complex nature of the “doubt” emotion and the importance of considering multiple modalities for accurate recognition and reasoning.
@@
Table 14: Example of the emotion category “contempt”. The table showcases a multimodal data sample corresponding to the “contempt” emotion, encompassing the visual expression description, audio tone description, and textual content. This example illustrates the nuanced and challenging characteristics of the “contempt” category, emphasizing the need for a comprehensive multimodal approach to capture its subtleties.
@@ B.3.1 Multi-Task Learning Approach
One of the core aspects of the Emotion-LLaMA model is its ability to simultaneously learn and perform multiple tasks related to emotion understanding. Specifically, the model is trained using a multi-task learning approach, where the emotion recognition and emotion reasoning tasks are learned in parallel. In the emotion recognition task, the model learns to predict the appropriate emotion label for a given multimodal input, such as a video clip accompanied by audio and text. The model is trained to map the input features to one of the predefined emotion categories, such as happy, sad, angry, or neutral. On the other hand, the emotion reasoning task focuses on generating human-interpretable explanations for the predicted emotion labels. Given a multimodal input, the model learns to analyze the various cues and generate a natural language explanation that justifies the predicted emotion based on the available evidence. By training the model to perform both tasks simultaneously, we allow it to share representations and learn complementary skills across tasks. This multi-task learning approach has several benefits. First, it enables the model to develop a more comprehensive understanding of emotions by learning to recognize the emotional state and explain the reasoning behind it. Second, it encourages the model to learn more robust and generalizable representations by leveraging the commonalities and differences between the two tasks. During training, for each video sample input into the model, we randomly select either the emotion recognition or emotion reasoning task to enhance the model’s generalization and robustness.
@@ B.3.2 Coarse-to-Fine Training Strategy
To facilitate the learning of nuanced and detailed emotional representations, we employ a coarse-to-fine training strategy. This strategy involves two main stages: pre-training on coarse-grained annotations and fine-tuning on fine-grained annotations. In the pre-training stage, the model is trained on the coarse-grained annotations from the MERR dataset. These annotations provide a simple description of emotions, such as visual expression, audio tone. By training on these coarse-grained annotations, the model learns to capture the general emotional tone and develop an initial understanding of the emotional content in the multimodal data. After the pre-training stage, we proceed to the fine-tuning stage, where the model is exposed to the fine-grained annotations from the MERR dataset. These annotations offer more detailed and specific emotion descriptions, such as environmental information, body movements, and the emotional nuances inferred from tone and textual subtitles. By fine-tuning the model on these fine-grained descriptions, we enhance its emotional understanding and enable it to capture more nuanced expressions and variations in scenarios that closely resemble real-world contexts. The coarse-to-fine training strategy offers several advantages. First, it allows the model to gradually learn more complex and detailed emotional representations, starting from a solid foundation of general emotional understanding. Second, it helps to prevent overfitting and improves the model’s generalization ability by providing a hierarchy of emotional annotations to learn from. During the fine-tuning stage, we typically use a smaller learning rate and a shorter training duration compared to the pre-training stage. This allows the model to adapt its representations to the fine-grained labels without drastically changing the learned features from the pre-training stage.